# Computing Chebyshev polynomials using the complex Remez algorithm

### v2

@@
Abstract
We employ the generalized Remez algorithm, initially suggested by P. T. P. Tang, to perform an experimental study of Chebyshev polynomials in the complex plane. Our focus lies particularly on the examination of their norms and zeros. What sets our study apart is the breadth of examples considered, coupled with the fact that the degrees under investigation are substantially higher than those in previous studies where other methods have been applied. These computations of Chebyshev polynomials of high degrees reveal discernible patterns which allow for conjectures to be formulated based on abundant experimental evidence. The use of Tang’s algorithm allows for computations executed with precision, maintaining accuracy within quantifiable margins of error. Additionally, as a result of our experimental study, we propose what we believe to be a fundamental relationship between Chebyshev and Faber polynomials associated with a compact set.
1 Introduction
Let $\mathsf{E}$ be a compact subset of the complex plane $\mathbb{C}$. Our focus is directed towards monic polynomials that exhibit minimal deviation from zero over the set $\mathsf{E}$. In other words, for any given positive integer $n$ we want to find coefficients $a_{0}^{\ast},\dotsc,a_{n-1}^{\ast}$ satisfying
$\max_{z\in\mathsf{E}}\left|z^{n}+\sum_{k=0}^{n-1}a_{k}^{\ast}z^{k}\right|=\min_{a_{0},\dotsc,a_{n-1}\in\mathbb{C}}\max_{z\in\mathsf{E}}\left|z^{n}+\sum_{k=0}^{n-1}a_{k}z^{k}\right|.$ (1)
The existence of minimizing coefficients $a_{0}^{\ast},\dotsc,a_{n-1}^{\ast}$ is guaranteed through a compactness argument. However, such a minimizer does not need to be unique. If $\mathsf{E}$ is a finite points set consisting of $m<n$ points then there are an infinite number of different minimizing polynomials. This is the only exceptional case and the assumption that $\mathsf{E}$ consists of infinitely many points ensures the uniqueness of a monic minimizer of (1) for any $n$ which henceforth is denoted by $T_{n}^{\mathsf{E}}$. This is the so-called Chebyshev polynomial of degree $n$ corresponding to the set $\mathsf{E}$. For basic theory detailing the proofs of existence and uniqueness of Chebyshev polynomials we refer the reader to [2, 7, 9, 36, 25, 41].
@@ 1 Introduction
Historically the consideration of polynomial minimizers with respect to the maximum norm originate from the studies of P. L. Chebyshev who considered minimization on $[-1,1]$, see [6]. Chebyshev polynomials corresponding to real sets have been much better understood than the corresponding complex ones. The reason for this discrepancy in understanding can partially be attributed to the powerful alternation theorem which is valid for real Chebyshev polynomials, [7, p. 75]. For any compact set $\mathsf{E}\subset\mathbb{R}$ containing at least $n+1$ points the Chebyshev polynomial $T_{n}^{\mathsf{E}}$ is characterized by having an alternating set on $\mathsf{E}$ consisting of $n+1$ points. That is to say, there are points $x_{0}<x_{1}<\cdots<x_{n}$ all contained in $\mathsf{E}$ such that
$T_{n}^{\mathsf{E}}(x_{k})=(-1)^{n-k}\|T_{n}^{\mathsf{E}}\|_{\mathsf{E}}.$ (2)
@@
This alternating property, whose analogue can be shown for general real approximation tasks, constitutes the theoretical grounding for the classical Remez algorithm which is used to compute real-valued best approximations, see [7, 25, 34, 35].
1.1 Chebyshev polynomials in the complex plane
@@
Alternation fails to characterise Chebyshev polynomials for general complex sets $\mathsf{E}\subset\mathbb{C}$. Apart from the fact that the argument of a Chebyshev polynomial at an extremal point can be any angle, not just $k\pi$ with $k\in\mathbb{Z}$, the number of extremal points corresponding to $T_{n}^{\mathsf{E}}$ on $\mathsf{E}$ can vary greatly. While there are at least $n+1$ such extremal points on $\mathsf{E}$, see e.g. [41, Theorem 1, p. 446], there is no upper bound on the number of extremal points. Indeed, as the example $T_{n}^{\mathbb{T}}(z)=z^{n}$ shows, the entire sets may consist of extremal points of the Chebyshev polynomial.
@@
One approach to studying Chebyshev polynomials in the complex plane comes from the fruitful interplay between approximation theory and potential theory. For this reason, we recall that to any compact set $\mathsf{E}\subset\mathbb{C}$ we can associate a quantity referred to as the logarithmic capacity, denoted $\operatorname{Cap}(\mathsf{E})$, see [33, §5.1].
In [44] Szegő proved that
$\|T_{n}^{\mathsf{E}}\|_{\mathsf{E}}\geq\operatorname{Cap}(\mathsf{E})^{n}.$ (3)
@@
A recent proof of this fundamental inequality can be found in [33, Theorem 5.5.4]. Since the capacity and radius of a disk coincides this provides an easy way of seeing that $T_{n}^{\mathbb{T}}(z)=z^{n}$. However, this powerful inequality can be used to draw further conclusions. If $P(z)=a_{m}z^{m}+a_{m-1}z^{m-1}+\cdots+a_{0}$ is a polynomial of exact degree $m$, then [33, Theorem 5.2.5] says that
$\operatorname{Cap}\Big{(}P^{-1}(\mathsf{E})\Big{)}=\left(\frac{\operatorname{Cap}(\mathsf{E})}{|a_{m}|}\right)^{1/m}.$ (4)
@@
Let $P$ be a monic polynomial of degree $m$ and $\mathsf{E}_{P}=\{z:|P(z)|\leq r\}$ a filled-in lemniscate. We gather as a consequence of (3), (4) and the uniqueness of Chebyshev polynomials that
$T_{nm}^{\mathsf{E}_{P}}(z)=P(z)^{n}.$ (5)
@@
This example, whose origins can be traced back to Faber [14], constitutes one of the few cases where the Chebyshev polynomials are explicitly determined for certain degrees. In general, for a given compact set $\mathsf{E}$, it is rarely the case that the Chebyshev polynomials $T_{n}^{\mathsf{E}}$ have known representations. For instance, the Chebyshev polynomials corresponding to $\{z:|P(z)|\leq r\}$ of degrees other than multiples of $\deg(P)$ remain unknown in the general case.
Chebyshev polynomials appear in various applications. The classical Chebyshev polynomials which are minimal on intervals are fundamental for numerical analysis and approximation theory. This is, to a large extent, due to their relation with Fourier analysis. Chebyshev polynomials on unions of intervals further appear as discriminants corresponding to Jacobi matrices which in turn are related to periodic Schrödinger operators, see [9, §2]. The generalization of Chebyshev polynomials to complex sets can also be motivated by applicability. To name an example, it is explained in [19] how matrix valued Chebyshev polynomials have applications to Krylov subspace iterations methods such as the Arnoldi iteration which is used to estimate eigenvalues of matrices. Such potential links are further considered in [48, 15]. If the matrix in question is normal then the matrix valued Chebyshev polynomials coincide with Chebyshev polynomials relative to the spectrum of the matrix in question. On a related note, residual matrix valued Chebyshev polynomials appear when estimating convergence of the GMRES algorithm, see [19]. Residual Chebyshev polynomials are also minimizers of the supremum norm on a compact set but instead of being monic they are normalized to attain the value $1$ at some specified point. This modification gives rise to differences but many properties are shared. For theoretical aspects of such polynomials see [12]. We believe that these examples serves to indicate that the determination of Chebyshev polynomials is interesting for a variety of different reasons and not limited to understanding fundamental properties of approximation theory.
1.2 Two different approaches
@@
To remedy the fact that Chebyshev polynomials typically are inexplicit, one common approach to understanding their asymptotic behavior is to compare them to explicit classes of polynomials. One such class of polynomials are the Faber polynomials [14]. If $\mathsf{E}\subset\mathbb{C}$ is a simply connected compact set which consists of more than one point, there exists a conformal mapping $\Phi:\mathbb{C}\setminus\mathsf{E}\rightarrow\mathbb{C}\setminus\overline{\mathbb{D}}$ of the form
$\Phi(z)=\operatorname{Cap}(\mathsf{E})^{-1}z+a_{0}+a_{-1}z^{-1}+\cdots$ (6)
@@
see [41, Chapter 2]. The Faber polynomial of degree $n$, denoted $F_{n}^{\mathsf{E}}$, is the monic polynomial of degree $n$ defined by the equation
$\Big{(}\operatorname{Cap}(\mathsf{E})\Phi(z)\Big{)}^{n}=F_{n}^{\mathsf{E}}(z)+O(z^{-1}),\quad z\rightarrow\infty.$ (7)
@@
In certain rare cases the Chebyshev polynomials and Faber polynomials corresponding to a set coincide [14]. If $\mathsf{E}$ is the closure of an analytic Jordan domain the Faber polynomials become well-suited trial polynomials for studying Chebyshev polynomials. In this case they satisfy
$\|F_{n}^{\mathsf{E}}\|_{\mathsf{E}}=\operatorname{Cap}(\mathsf{E})^{n}\Big{(}1+O(r^{n})\Big{)}$ (8)
@@
for some $10<r<$ 1 depending on $\mathsf{E}$, see e.g. [14, 51]. This implies that the sequence $\{F_{n}^{\mathsf{E}}\}$ asymptotically saturates (3) when the bounding curve of $\mathsf{E}$ is analytic. As a consequence, the so-called Widom factors introduced in [18] as
$\mathcal{W}_{n}(\mathsf{E}):=\frac{\|T_{n}^{\mathsf{E}}\|_{\mathsf{E}}}{\operatorname{Cap}(\mathsf{E})^{n}}$ (9)
@@
converge to $1$ as $n\rightarrow\infty$ in this particular case. From (3) we see that this is optimal. Much of the research into Chebyshev polynomials is directed to understanding the asymptotic behavior of $\mathcal{W}_{n}(\mathsf{E})$. In [51] and later [42] conditions to guarantee that $\mathcal{W}_{n}(\mathsf{E})\rightarrow 1$ as $n\rightarrow\infty$ were relaxed by means of comparison with Faber polynomials. If $\mathsf{E}$ is the closure of a Jordan domain with $C^{1+\alpha}$ boundary then it follows from [43, Theorem 2, p.68] that
$\mathcal{W}_{n}(\mathsf{E})\leq\frac{\|F_{n}\|}{\operatorname{Cap}(\mathsf{E})^{n}}=1+O\left(\frac{\log n}{n^{\alpha}}\right).$ (10)
@@
It is an open question if the conditions concerning the regularity of the boundary can be further relaxed while still guaranteeing that the corresponding Widom factors converge to the theoretical minimal value. For instance, if $\mathsf{E}$ is the closure of a Jordan domain such that the bounding curve is piecewise analytic but contains corner points can we still conclude that
@@
Using Faber polynomials it can be shown that if $\mathsf{E}$ is a convex set then $\mathcal{W}_{n}(\mathsf{E})\leq 2$ for all $n$, see [28, Theorem 2] and more recently [1, Theorem 4.1]. In [4] a completely different class of trial polynomials were used to prove that the sequence $\{\mathcal{W}_{n}(\mathsf{E})\}$ remains bounded if $\mathsf{E}$ is the closure of a quasi-disk. For examples illustrating the close interplay between Faber polynomials and Chebyshev polynomials, we refer the reader to [40, 51]. In Section 3 we will explore a possible relation between the Chebyshev and Faber polynomials that have been observed numerically. Loosely formulated this entails that Chebyshev polynomials approach Faber polynomials for a fixed degree along certain curves related to the conformal map of the set in question.
@@
Besides understanding the norm behavior, another point of interest is understanding how the geometry of a set affects the zero distributions of the corresponding Chebyshev polynomials. Given a polynomial $P$, let $\nu(P)$ denote the normalized zero counting measure of $P$. That is,
$\nu(P)=\frac{1}{\deg(P)}\sum_{j=1}^{\deg(P)}\delta_{z_{j}}$
@@
where $\delta_{z}$ is the Dirac delta measure at $z$ and $\{z_{j}\}$ denotes the zeros of $P$ counting multiplicity. Given a compact set $\mathsf{E}$, a typical quantitative way of describing the asymptotical distribution of the zeros of $T_{n}^{\mathsf{E}}$ is by determining weak-star limits of the sequence of measures $\{\nu(T_{n}^{\mathsf{E}})\}$. As it turns out, such weak-star limits are closely related to the potential theoretic concept of equilibrium measure. We therefore introduce the notation $\mu_{\mathsf{E}}$ to denote the equilibrium measure corresponding to a compact set $\mathsf{E}$, see [33, §3.3]. Given a sequence of degrees $\{n_{k}\}$ [3, Theorem 2.1.7] says that if
$\lim_{k\rightarrow\infty}\nu(T_{n_{k}}^{\mathsf{E}})(M)=0$ (11)
@@
for every compact set $M$ in the interior of $\mathsf{E}$ then $\nu(T_{n_{k}}^{\mathsf{E}})\xrightarrow{\ast}\mu_{\mathsf{E}}$ as $n_{k}\rightarrow\infty$. Loosely formulated, if “almost all” of the zeros of $T_{n_{k}}^{\mathsf{E}}$ approach the boundary then they distribute according to equilibrium measure. In particular, if $\mathsf{E}$ has empty interior then $\nu(T_{n}^{\mathsf{E}})\xrightarrow{\ast}\mu_{\mathsf{E}}$ as $n\rightarrow\infty$. It is shown in [40] that the zeros of $T_{n}^{\mathsf{E}}$ when $\mathsf{E}$ is the closure of a Jordan domain, stay away from the boundary precisely when the bounding curve is analytic. It therefore follows that if $\mathsf{E}$ is the closure of a Jordan domain whose boundary contains a corner then the zeros of $T_{n}^{\mathsf{E}}$ will approach the boundary in some fashion. The question we want to investigate is if we can discern that (11) should hold for such sets.
@@
It could be argued that in order to study Chebyshev polynomials there are two available approaches. One alternative is to try to compare Chebyshev polynomials with other classes of polynomials which are candidates to provide small maximum norms such as the Faber polynomials. The other approach to studying Chebyshev polynomials – which will be the main focus of this article – is to consider computing these polynomials. In our case these computations will be performed using numerical approximations. Such considerations are somewhat scarce in the literature although examples exist which rely on other methods than the ones presented here. See for instance [20, 32, 47, 48]. In this article we will discuss and apply an algorithm suggested by P. T. P. Tang that was presented in his Ph.D thesis [45] and further developed by B. Fischer and J. Modersitzki in [16]. More specifically we will compute Chebyshev polynomials corresponding to a wide variety of compact sets in the complex plane. Doing so, it will become apparent that certain hypothesis can be made plausible using numerical computations. See [16, 27, 46] for further developments of this algorithm.
1.3 Outline

### v3

@@
Abstract.
We employ the generalized Remez algorithm, initially suggested by P. T. P. Tang, to perform an experimental study of Chebyshev polynomials in the complex plane. Our focus lies particularly on the examination of their norms and zeros. What sets our study apart is the breadth of examples considered, coupled with the fact that the degrees under investigation are substantially higher than those in previous studies where other methods have been applied. These computations of Chebyshev polynomials of high degrees reveal discernible, repeating patterns, which indicate a typical behavior of Chebyshev polynomials in a general setting. The use of Tang’s algorithm allows for computations executed with precision, maintaining accuracy within quantifiable margins of error. Additionally, as a result of our experimental study, we propose what we believe to be a fundamental relationship between Chebyshev and Faber polynomials associated with a compact set.
Key words and phrases:
Chebyshev polynomials, Remez algorithm, Widom factors, Zeros of polynomials.
2020 Mathematics Subject Classification:
41A50, 65E05, 30C10
1. Introduction
Let $\mathsf{E}$ be a compact subset of the complex plane $\mathbb{C}$. Our focus is directed towards monic polynomials that exhibit minimal deviation from zero over the set $\mathsf{E}$. In other words, for any given positive integer $n$ we want to find the coefficients $a_{0}^{\ast},\dotsc,a_{n-1}^{\ast}$ satisfying
(1) $\max_{z\in\mathsf{E}}\left|z^{n}+\sum_{k=0}^{n-1}a_{k}^{\ast}z^{k}\right|=\min_{a_{0},\dotsc,a_{n-1}\in\mathbb{C}}\max_{z\in\mathsf{E}}\left|z^{n}+\sum_{k=0}^{n-1}a_{k}z^{k}\right|.$
The existence of minimizing coefficients $a_{0}^{\ast},\dotsc,a_{n-1}^{\ast}$ is guaranteed through a compactness argument. However, such a minimizer does not need to be unique. If $\mathsf{E}$ is finite and consists of $m<n$ points, then there are infinitely many different minimizing polynomials. This is the only exceptional case, the assumption that $\mathsf{E}$ consists of infinitely many points ensures the uniqueness of a monic $n$th degree minimizer of (1) for any $n\in\mathbb{N}$. This is the so-called Chebyshev polynomial of degree $n$ corresponding to the set $\mathsf{E}$ which we denote by $T_{n}^{\mathsf{E}}$. For basic theory detailing the proofs of existence and uniqueness of Chebyshev polynomials we refer the reader to [2, 8, 10, 36, 41, 27, 46].
@@ 1. Introduction
Historically, the consideration of polynomial minimizers with respect to the maximum norm originates from the studies of P. L. Chebyshev who considered minimization on $[-1,1]$, see [7]. Chebyshev polynomials corresponding to real sets have been much better understood than the corresponding complex ones. The reason for this discrepancy in understanding can partially be attributed to the powerful alternation theorem, which is valid for real Chebyshev polynomials, [8, p. 75]. For any compact set $\mathsf{E}\subset\mathbb{R}$ containing at least $n+1$ points, the Chebyshev polynomial $T_{n}^{\mathsf{E}}$ is characterized by having an alternating set on $\mathsf{E}$ consisting of $n+1$ points. That is to say, there are points $x_{0}<x_{1}<\cdots<x_{n}$ all contained in $\mathsf{E}$ such that
(2) $T_{n}^{\mathsf{E}}(x_{k})=(-1)^{n-k}\|T_{n}^{\mathsf{E}}\|_{\mathsf{E}}.$
@@
This alternating property, whose analogue can be shown for any real best approximations, constitutes the theoretical grounding for the classical Remez algorithm which is used to compute real-valued best approximations, see [8, 27, 39, 40].
1.1. Chebyshev polynomials in the complex plane
@@
Crucially, the alternation property fails to characterize Chebyshev polynomials for general complex sets $\mathsf{E}\subset\mathbb{C}$. Apart from the fact that the argument of a Chebyshev polynomial at an extremal point can be any angle, not just $k\pi$ with $k\in\mathbb{Z}$, the number of extremal points corresponding to $T_{n}^{\mathsf{E}}$ on $\mathsf{E}$ can vary greatly. While there are at least $n+1$ such extremal points on $\mathsf{E}$, see e.g. [46, Theorem 1, p. 446], there is no upper bound on the number of extremal points. Indeed, as the example $T_{n}^{\mathbb{T}}(z)=z^{n}$ shows, the entire sets may consist of extremal points of the Chebyshev polynomial.
@@
One approach to studying Chebyshev polynomials in the complex plane comes from the fruitful interplay between approximation theory and potential theory. We recall that the logarithmic capacity is a quantity associated with any compact set in the complex plane, see [38, §5.1]. Denoting it with $\operatorname{Cap}$, Szegő [49] proved that
(3) $\|T_{n}^{\mathsf{E}}\|_{\mathsf{E}}\geq\operatorname{Cap}(\mathsf{E})^{n}.$
@@
A recent proof of this fundamental inequality can be found in [38, Theorem 5.5.4]. Since the capacity and radius of a disk coincide, this provides an easy way of seeing that $T_{n}^{\mathbb{T}}(z)=z^{n}$. If $P(z)=a_{m}z^{m}+a_{m-1}z^{m-1}+\cdots+a_{0}$ is a polynomial of exact degree $m$, then [38, Theorem 5.2.5] says that
(4) $\operatorname{Cap}\Big{(}P^{-1}(\mathsf{E})\Big{)}=\left(\frac{\operatorname{Cap}(\mathsf{E})}{|a_{m}|}\right)^{1/m}.$
@@
If we additionally assume that $P$ is a monic polynomial of degree $m$ and define the filled-in lemniscate $\mathsf{E}_{P}=\{z:|P(z)|\leq r\}$, then we gather as a consequence of (3), (4) and the uniqueness of Chebyshev polynomials that
(5) $T_{nm}^{\mathsf{E}_{P}}(z)=P(z)^{n}.$
@@
This example, whose origin can be traced back to Faber [15], constitutes one of the few cases where the Chebyshev polynomials are determined for certain degrees. In general, for a given compact set $\mathsf{E}$, it is rarely the case that the Chebyshev polynomials $T_{n}^{\mathsf{E}}$ admit explicit representations. For instance, the Chebyshev polynomials corresponding to $\{z:|P(z)|\leq r\}$ of degrees other than multiples of $\deg(P)$, remain unknown in the general case.
Chebyshev polynomials appear in various applications. The classical Chebyshev polynomials, which minimize the supremum norm on the interval $[-1,1]$, are fundamental for numerical analysis and approximation theory. This is, to a large extent, due to their relation with Fourier analysis. Chebyshev polynomials on unions of intervals further appear as discriminants corresponding to Jacobi matrices, which in turn are related to periodic Schrödinger operators, see [10, §2].
The extension of Chebyshev polynomials to complex sets can also be motivated by their applicability. For example, [21] explains how matrix-valued Chebyshev polynomials have applications to iterative Krylov subspace methods, such as the Arnoldi iteration, which is used to estimate eigenvalues of matrices. Such potential links are further considered in [53, 16]. If the matrix in question is normal then the matrix-valued Chebyshev polynomials coincide with the Chebyshev polynomials on the spectrum of the corresponding matrix.
Closely related, residual Chebyshev polynomials are also minimizers of the supremum norm on a compact set, but instead of being monic they are normalized to attain the value $1$ at some specified point. When the reference point is at infinity this normalization should be understood as fixing the leading coefficient to be $1$ and hence the Chebyshev polynomials emerge. This modification gives rise to differences but many properties are shared. Residual matrix-valued Chebyshev polynomials appear while estimating the convergence rate of the GMRES algorithm, see [21]. For theoretical aspects of residual Chebyshev polynomials see [13]. While we will focus on Chebyshev polynomials, the numerical methods presented here can be applied to the residual case as well.
We believe these examples indicate that determining Chebyshev polynomials is interesting for various reasons and not limited to understanding fundamental properties of approximation theory.
1.2. Two different approaches
@@
To remedy the fact that Chebyshev polynomials typically lack explicit formulas, one common approach to understand their asymptotic behavior is to compare them to other classes of polynomials. One such class of polynomials is the Faber polynomials [15]. If $\mathsf{E}\subset\mathbb{C}$ is a simply connected compact set which consists of more than one point, then there exists a conformal mapping $\Phi:\mathbb{C}\setminus\mathsf{E}\rightarrow\mathbb{C}\setminus\overline{\mathbb{D}}$ of the form
(6) $\Phi(z)=\operatorname{Cap}(\mathsf{E})^{-1}z+a_{0}+a_{-1}z^{-1}+\cdots,$
@@
see [46, Chapter 2]. The (normalized) $n$th Faber polynomial, denoted $F_{n}^{\mathsf{E}}$, is the monic polynomial of degree $n$ defined by the equation
(7) $\Big{(}\operatorname{Cap}(\mathsf{E})\Phi(z)\Big{)}^{n}=F_{n}^{\mathsf{E}}(z)+O(z^{-1}),\quad z\rightarrow\infty.$
@@
In certain rare cases, the Chebyshev polynomials and Faber polynomials corresponding to a set coincide [15]. In other cases, the Faber polynomials provide near minimal upper bounds. If $\mathsf{E}$ is the closure of an analytic Jordan domain then
(8) $\|F_{n}^{\mathsf{E}}\|_{\mathsf{E}}=\operatorname{Cap}(\mathsf{E})^{n}\Big{(}1+O(r^{n})\Big{)}$
@@
for some $10<r<$ 1, see e.g. [15, 57]. This implies that the sequence $\{F_{n}^{\mathsf{E}}\}$ asymptotically saturates (3) in this case and
$\frac{\|T_{n}^{\mathsf{E}}\|_{\mathsf{E}}}{\operatorname{Cap}(\mathsf{E})^{n}}\rightarrow 1,\quad\text{as}\quad n\rightarrow\infty$
which is optimal as seen from (3). The Widom factor was introduced in [20] and is defined as
(9) $\mathcal{W}_{n}(\mathsf{E})\coloneqq\frac{\|T_{n}^{\mathsf{E}}\|_{\mathsf{E}}}{\operatorname{Cap}(\mathsf{E})^{n}}.$
@@
Much of the research into Chebyshev polynomials is directed to understanding the asymptotic behavior of $\mathcal{W}_{n}(\mathsf{E})$. In [57] and later [47] conditions to guarantee that $\mathcal{W}_{n}(\mathsf{E})\rightarrow 1$ as $n\rightarrow\infty$ were relaxed by means of comparison with Faber polynomials. If $\mathsf{E}$ is the closure of a Jordan domain with $C^{1+\alpha}$ boundary then it follows from [48, Theorem 2, p.68] that
(10) $\mathcal{W}_{n}(\mathsf{E})\leq\frac{\|F_{n}\|}{\operatorname{Cap}(\mathsf{E})^{n}}=1+O\left(\frac{\log n}{n^{\alpha}}\right).$
@@
It is an open question if the conditions concerning the regularity of the boundary can be further relaxed while still guaranteeing that the corresponding Widom factors converge to the theoretical minimal value. For instance, if $\mathsf{E}$ is the closure of a Jordan domain such that the bounding curve is piecewise-analytic but contains corner points can we still conclude that
@@
Through a comparison with Faber polynomials it can be shown that if $\mathsf{E}$ is a convex set then $\mathcal{W}_{n}(\mathsf{E})\leq 2$ for all $n$. A proof of this fact using an upper bound on the norm of the corresponding Faber polynomials deduced from [30, Theorem 2] can be found in [42, Theorem 10]. A strengthened form of the estimate is given in [1, Theorem 4.1].
@@
Faber polynomials are not the only polynomials useful for studying Chebyshev polynomials. A completely different class of trial polynomials were used in [5] to prove that the sequence $\{\mathcal{W}_{n}(\mathsf{E})\}$ remains bounded if $\mathsf{E}$ is the closure of a quasi-disk. For examples illustrating the close interplay between Faber polynomials and Chebyshev polynomials, we refer the reader to [45, 57]. In Section 3 we will explore a possible relation between the Chebyshev and Faber polynomials that has been observed numerically. Loosely formulated this entails that Chebyshev polynomials approach Faber polynomials for a fixed degree along certain curves related to the conformal map of the set in question.
@@
Besides understanding the norm behavior, another point of interest lies in understanding how the geometry of a set affects the zero distributions of the corresponding Chebyshev polynomials. Given a polynomial $P$, let $\nu(P)$ denote the normalized zero-counting measure of $P$. That is,
$\nu(P)=\frac{1}{\deg(P)}\sum_{j=1}^{\deg(P)}\delta_{z_{j}},$
@@
where $\delta_{z}$ is the Dirac delta measure at $z$ and $\{z_{j}\}$ denotes the zeros of $P$ counting multiplicity. Given a compact set $\mathsf{E}$, a typical quantitative way of describing the asymptotical distribution of the zeros of $T_{n}^{\mathsf{E}}$ is by determining weak-star limits of the sequence of measures $\{\nu(T_{n}^{\mathsf{E}})\}$. As it turns out, such weak-star limits are closely related to the potential theoretic concept of the equilibrium measure. We therefore introduce the notation $\mu_{\mathsf{E}}$ to denote the equilibrium measure corresponding to a compact set $\mathsf{E}$, see [38, §3.3]. Given a sequence of degrees $\{n_{k}\}$, [4, Theorem 2.1.7] says that if
(11) $\lim_{k\rightarrow\infty}\nu(T_{n_{k}}^{\mathsf{E}})(M)=0$
@@
for every compact set $M$ in the interior of $\mathsf{E}$ then $\nu(T_{n_{k}}^{\mathsf{E}})\xrightarrow{\ast}\mu_{\mathsf{E}}$ as $n_{k}\rightarrow\infty$. Loosely formulated, if “almost all” of the zeros of $T_{n_{k}}^{\mathsf{E}}$ approach the boundary then they distribute according to equilibrium measure. In particular, if $\mathsf{E}$ has empty interior then $\nu(T_{n}^{\mathsf{E}})\xrightarrow{\ast}\mu_{\mathsf{E}}$ as $n\rightarrow\infty$. It is shown in [45] that the zeros of $T_{n}^{\mathsf{E}}$, when $\mathsf{E}$ is the closure of a Jordan domain, stay away from the boundary precisely when the bounding curve is analytic. It therefore follows that if $\mathsf{E}$ is the closure of a Jordan domain whose boundary contains a corner, then the zeros of $T_{n}^{\mathsf{E}}$ will approach the boundary in some fashion. The question we want to investigate is if we can discern that (11) is plausible for such sets.
@@
It could be argued that in order to study Chebyshev polynomials there are two available approaches. One way is to try to compare Chebyshev polynomials with other classes of polynomials which are candidates to provide small maximum norms such as the Faber polynomials. The other approach to studying Chebyshev polynomials – which will be the main focus of this article – is to consider computing these polynomials. In our case, these computations will be performed using numerical approximations. Such considerations are somewhat scarce in the literature, although examples exist which rely on other methods than the ones presented here. See for instance [18, 22, 37, 52, 53]. In this article we will discuss and apply an algorithm suggested by P. T. P. Tang that was presented in his PhD thesis [50] and further developed by B. Fischer and J. Modersitzki in [17]. More specifically, we will compute Chebyshev polynomials corresponding to a wide variety of compact sets in the complex plane. Doing so, it will become apparent that certain hypothesis can be made plausible using numerical computations. See [17, 29, 51] for further developments of this algorithm.
Another, well-known algorithm which could be used for the purpose of computing Chebyshev polynomials is the Lawson algorithm and its extensions, see [33]. One big difference between Tang’s algorithm and the Lawson algorithm is that the Lawson algorithm computes the best approximation on a discrete subset. A detailed comparison between the performance of different algorithms that can be used to compute best approximations is conducted in [50].
For examples of studies of orthogonal polynomials based on experimental approaches, see [3, 31].
1.3. Outline